# Heavy quarkonium spectral function in an anisotropic background

### v1

@@ Abstract
In this paper, we use a five-dimensional Einstein-dilaton-two-Maxwell holographic QCD model to investigate the dissociation effects of $J/\Psi$ and $\Upsilon(1S)$ states in an anisotropic medium by calculating their spectral functions. The anisotropic background is characterized by an anisotropy parameter, a nontrivial warp factor, a dilaton field, and distinct components of two Maxwell fields. First, we present the holographic quarkonium masses at zero temperature via Physics-Informed Neural Networks. Then, at finite temperature, we derive the spectral functions, representing heavy vector mesons as peaks, and observe that with increasing anisotropy, temperature, chemical potential, and warp factor, the peak height diminishes while its width expands, indicating an accelerated dissociation process. Additionally, the results indicate the anisotropy induces a stronger dissociation effect in the direction parallel to the polarization compared to the perpendicular, revealing the anisotropy’s directional influence.
@@ II Anisotropic Holographic Model
The chemical potential and anisotropy are introduced by two Maxwell fields $F_{(1)}$ and $F_{(2)}$, with field strength tensors $F_{\mu\nu}^{(1)}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $F_{\mu\nu}^{(2)}=qdy^{1}\wedge dy^{2}$, respectively. The dilaton field $\phi$, with potential $V(\phi)$, couples to these two Maxwell fields through the gauge kinetic functions $f_{1}(\phi)$ and $f_{2}(\phi)$. To describe the anisotropic background holographically, we adopt the black brane metric ansatz in the following form aref :
@@
Here, $b(z)=e^{cz^{2}/2}$ represents the AdS deformation factor, with $c$ signifying the deviation from conformality. $g(z)$ is the blackening function, and $L$ is the characteristic length scale (set to one for convenience). A non-zero time component, $A_{t}(0)=\mu$, of the first Maxwell field $F_{(1)}$, serves as the chemical potential in the dual system. An intriguing aspect of this model is the anisotropy parameter $\nu$, which serves as a measure of the degree of anisotropy in the system. The ansatz exhibits rotational invariance in the $x_{2}$$x_{3}$ plane with the $x_{1}$ as the direction of anisotropy, breaking isotropy while preserving translation and boost invariances. By solving the equation of motion derived from the aforementioned action, the functioncan $g(z)$ be determined as
@@ IV RESULTS AND DISCUSSION
Physics-Informed Neural Networks (PINNs) is an innovative approach that embeds physical laws directly into neural network training by treating these laws as regularization terms or constraints within the loss function. It presents a powerful methodology across diverse domains, including fluid dynamics and quantum mechanics, offering some advantages over conventional numerical methods, such as mitigating the curse of dimensionality, minimizing numerical error accumulation, bolstering computational efficiency through inherent parallelizability, and obviating the need for tailor-made numerical approaches for each individual case pgr ; mra ; chenx . PINNs offer a simple implementation, requiring minimal code in any machine learning framework with automatic differentiation, and can process differential equations without specific requirements, provided their solutions are finite and regular Teuko . Jin, Mattheakis, and Protopapas have recently developed a PINNs algorithm that simultaneously learns eigenvalues and eigenfunctions through a scanning mechanism, effectively solving the quantum eigenvalue problem for finite wells, harmonic oscillators, and hydrogen atom systems by progressively identifying multiple eigenvalues in a single training session jin1 ; jin2 .
In this study, we present an extension of the JMP algorithm to tackle more complex potential function scenarios, generalizing its application to solve the mass of quarkonium. Further details on the PINNs method used in our calculations are provided in Appendix A. Figures 1 and 2 depict the evolution of the loss function and show the successful recognition of the four lowest eigenvalues of charmonium and bottomonium by our neural network.
@@
The neural network identifies eigenvalues by searching for plateaus, which correspond to distinct depressions in the loss function, with a jump caused by introducing $L_{drive}$ to encourage finding the next eigenvalue. The summarized results of the neural network for charmonium and bottomonium are shown in tables 1 and 2, respectively, with the experimental data read from PDG pdg for comparison.
@@ V Acknowledgments
We thank Yan-Qing Zhao for the fruitful discussions. This work is supported in part by the National Key Research and Development Program of China under Contract No. 2022YFA1604900. This work is also partly supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12275104, No. 11890711, No. 11890710, and No. 11735007.
@@
Figure 12 illustrates the algorithm’s framework, designed to address equation (23) and solve the eigenvalue by learning from their own predictions based on the differential equations and boundary conditions without requiring any labeled datasets jin1 ; jin2 .
Figure 12: The structure of the neural networks, implemented utilizing the PyTorch framework.
[…]
$\begin{gathered}f(x,\lambda)=f_{b}+g(x)N(x,\lambda),\\ g(x)=(1-e^{-(x-x_{L})})(1-e^{-(x-x_{R})}),\end{gathered}$ (24)
where the Dirichlet boundary conditions at the left and right boundaries are $x_{L}$  and $x_{R}$, respectively, while $f_{b}$ is the boundary value. To address equation (23), the network is optimized through the minimization of a loss function $L$
$\begin{gathered}L=\left\langle(\mathcal{L}f(x,\lambda)-\lambda f(x,\lambda))^{2}\right\rangle_{x}+L_{reg},\\ L_{reg}=L_{f}+L_{\lambda}+L_{drive},\end{gathered}$ (25)
where $\mathcal{L}$ represents the differential operator in (23), $L_{f}=\frac{1}{{f(x,\lambda)}^{2}}$, $L_{\lambda}=\frac{1}{\lambda^{2}}$, $L_{drive}=e^{-\lambda+c}$. $L_{f}$ and $L_{\lambda}$ are employed to prevent a trivial result, and $L_{drive}$ motivates the network to seek higher eigenvalues by incrementing the parameter $c$ during the training. In our investigation, we used PyTorch to implement a fully-connected neural network with two hidden layers, each comprising $100$ neurons, optimized using the Adam algorithm with a learning rate of $8\times 10^{-3}$ and a sine activation function, with the training set constructed by $2000$ random points within the $[x_{L},x_{R}]$ interval.

### v3

@@ Abstract
In this paper, we use a five-dimensional Einstein-dilaton-two-Maxwell holographic QCD model to investigate the dissociation effects of $J/\Psi$ and $\Upsilon(1S)$ states in an anisotropic medium by calculating their spectral functions. First, we present the holographic quarkonium masses at zero temperature via Physics-Informed Neural Networks. Then, at finite temperature, we derive the spectral functions, representing heavy vector mesons as peaks, and observe that with increasing anisotropy, temperature, chemical potential, and warp factor, the peak height diminishes while its width expands, indicating an accelerated dissociation process. Additionally, the results indicate the anisotropy induces a stronger dissociation effect in the direction parallel to the polarization compared to the perpendicular, revealing the anisotropy’s directional influence.
@@ II Anisotropic Holographic Model
The chemical potential and anisotropy are introduced by two Maxwell fields $F_{(1)}$ and $F_{(2)}$, with field strength tensors $F_{\mu\nu}^{(1)}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $F_{\mu\nu}^{(2)}=qdy^{1}\wedge dy^{2}$, respectively. The dilaton field $\phi$, with potential $V(\phi)$, couples to these two Maxwell fields through the gauge kinetic functions $f_{1}(\phi)$ and $f_{2}(\phi)$. To describe the anisotropic background holographically, we adopt the black brane metric ansatz in the following form aref
@@
Here, $b(z)=e^{cz^{2}/2}$ represents the AdS deformation factor, with $c$ signifying the deviation from conformality. $g(z)$ is the blackening function, and $L$ is the characteristic length scale (set to one for convenience). A non-zero time component, $A_{t}(0)=\mu$, of the first Maxwell field $F_{(1)}$, serves as the chemical potential in the dual system. By solving the equation of motion derived from the aforementioned action, the functioncan $g(z)$ be determined as
@@ IV RESULTS AND DISCUSSION
Physics-Informed Neural Networks (PINNs) is an innovative approach that embeds physical laws directly into neural network training by treating these laws as regularization terms or constraints within the loss function. It presents a powerful methodology across diverse domains, including fluid dynamics and quantum mechanics, offering some advantages over conventional numerical methods, such as mitigating the curse of dimensionality, minimizing numerical error accumulation, and processing differential equations without specific requirements, provided their solutions are finite and regular pgr ; mra ; chenx ; Teuko . Jin, Mattheakis, and Protopapas have recently developed a PINNs algorithm that simultaneously learns eigenvalues and eigenfunctions through a scanning mechanism, effectively solving the quantum eigenvalue problem for finite wells, harmonic oscillators, and hydrogen atom systems by progressively identifying multiple eigenvalues in a single training session jin1 ; jin2 .
In this study, we present an extension of the JMP algorithm to tackle more complex potential function scenarios, generalizing its application to solve the mass of quarkonium. Further details on the JMP algorithm used in our calculations are provided in Appendix A. Figures 1 and 2 depict the evolution of the loss function and show the successful recognition of the four lowest eigenvalues of charmonium and bottomonium by our neural network.
Refer to captionRefer to caption
@@
The neural network identifies eigenvalues by searching for plateaus, which correspond to distinct depressions in the loss function, with a jump caused by changing $c$ to encourage finding the next eigenvalue. The summarized results of the neural network for charmonium and bottomonium are shown in tables 1 and 2, respectively, with the experimental data read from PDG pdg for comparison.
@@ V Acknowledgments
We thank Hai-Cang Ren and Yan-Qing Zhao for the fruitful discussions. This work is supported in part by the National Key Research and Development Program of China under Contract No. 2022YFA1604900. This work is also partly supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12275104, No. 11890711, No. 11890710, and No. 11735007.
@@
Figure 12 illustrates the framework of the JMP algorithm, which solves the eigenvalue problem by learning from its own predictions jin1 ; jin2 .
Figure 12: The structure of the neural network.
The output of the neural network is denoted by $N(z,m_{n}^{2})$, and the predicted eigenfunction is given by
$\psi_{n}(z)=(1-e^{-(z-z_{L})})(1-e^{-(z-z_{R})})N(z,m_{n}^{2}),$ (24)
where the left and right boundaries are $z_{L}$  and $z_{R}$. To address equation (23), the network is optimized through the minimization of a loss function $L$
$\begin{gathered}L=\left\langle(-\psi_{n}^{\prime\prime}(z)+U(z)\psi_{n}(z)-m_{n}^{2}\psi_{n}(z))^{2}\right\rangle_{z}+L_{reg},\\ L_{reg}=\frac{1}{{\psi_{n}(z)}^{2}}+\frac{1}{(m_{n}^{2})^{2}}+e^{-m_{n}^{2}+c}.\end{gathered}$ (25)
In our investigation, we used PyTorch to implement a fully-connected neural network with two hidden layers, each comprising $100$ neurons, optimized using the Adam algorithm with a learning rate of $8\times 10^{-3}$ and sigmoid activation function, with the training set constructed by $2000$ random points within the $[z_{L},z_{R}]$ interval.